# A Hurewicz-type Theorem for the Dynamic Asymptotic Dimension with Applications to Coarse Geometry and Dynamics

### v4

@@ 1 Introduction
More recently, Hurewicz-type theorems were developed for Gromov’s asymptotic dimension in [asymptotichurewicztheorem] and later in [Brodskiy2006AHT]. This paper uses similar methods as in [Brodskiy2006AHT], together with some original ideas, to prove a Hurewicz-type theorem for the dynamic asymptotic dimension (DAD) of Guentner, Willett, and Yu. This dimension theory was introduced in [dasdimGWY] as a dynamical analogue of the asymptotic dimension. It is related to conditions used in work on the Farrell-Jones conjecture on manifold topology [bartels] [bartels_et_al], as well as to the homology of group actions [bonicke2021dynamic] and the nuclear dimension of transformation group $C^{*}$-algebras [dasdimGWY, Theorem 8.6].
@@
Suppose $1\to\Delta\to\Gamma\to\Lambda\to 1$ is an exact sequence of countable groups and that $\Gamma$ is residually finite. Suppose $(N_{i})$ is a countable collection of finite index normal subgroups of $\Gamma$ which separates $\Delta/(\cap_{i}(N_{i}\cap\Delta))$ from $\Gamma/(\cap_{i}N_{i})$ (see LABEL:separable_subgroup_definition). This collection induces collections of finite index normal subgroups and hence odometers of $\Delta$ and $\Lambda$. Then if $\text{DAD}_{free}(\Delta\curvearrowright\widehat{\Delta}_{(N_{i}\cap\Delta)})=m$ and $\text{DAD}_{free}(\Lambda\curvearrowright\widehat{\Lambda}_{(N_{i}/\Delta\cap N_{i})})=n$, $\text{DAD}_{free}(\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})})\leq m+n$.
This extension theorem can further be applied to coarse geometry, specifically the study of box spaces of groups. These are spaces constructed out of finite quotients of a group. Just as the asymptotic dimension of a group itself can have interesting group-theoretic applications (see for instance [MR2431020]), there are likewise many correspondences between geometric properties of box spaces and group-theoretic properties. For example, coarse equivalence of box spaces encodes virtual isomorphism of groups [MR4091056], and amenability of groups is equivalent to property A of their box spaces [Roe2003LecturesOC, Propsotion 11.39]. Similar results also exist for a-T-menablity [https://doi.org/10.1112/blms/bdt045].
Box spaces provided some of the first examples of spaces which do not coarsely embed in hilbert space [Roe2003LecturesOC, Remark 11.27] and it has been shown that a group with a finite dimensional box space must be amenable [Roe2003LecturesOC, Remark 11.38 and Proposition 11.39]. Similar to how actions by finite dimensional amenable groups are thought to be finite dimensional, it is thought that a partial converse to this latter fact should also hold; that is, an amenable group with finite asymptotic dimension ought to have finite dimensional box spaces. Notwithstanding, there has been similarly little progress on this aside from the case of virtually nilpotent groups [boxspacesDT].
@@
Let $\Gamma$ be a countable group and $(N_{i})$ a sequence of finite index subgroups of $\Gamma$. Suppose $\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})}$ is the action by left multiplication of $\Gamma$ on its profinite completion along $(N_{i})$. Then $\text{DAD}_{free}(\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})})=\text{asdim}\Box_{(N_{i})}\Gamma$.
@@
Suppose $1\to\Delta\to\Gamma\xrightarrow{\pi}\Lambda\to 1$ is an exact sequence of countable groups and $(N_{i})$ is a sequence of finite index normal subgroups of $\Gamma$ which separates $\Delta/(\cap_{i}(N_{i}\cap\Delta))$ from $\Gamma/(\cap_{i}N_{i})$. This sequence induces sequences and hence box spaces of $\Delta$ and $\Lambda$. Then if $\text{asdim}\Box_{(N_{i}\cap\Delta)}\Delta=m$ and $\text{asdim}\Box_{(\pi(N_{i}))}\Lambda=n$, $\text{asdim}\Box_{(N_{i})}\Gamma\leq m+n$.
We have therefore established a special case of the (unproved) statement [finnsell2015asymptotic, Theorem A]. Our extra assumptions are that $\Gamma$ is countable, that the $N_{i}$ are normal, and that $(N_{i})$ is a sequence separating $\Delta$ from $\Gamma$. In [finnsell2015asymptotic, Theorem A], the family of subgroups are only assumed to be semiconjugacy separating which is weaker than assuming $(N_{i})$ to be a filtration by normal subgroups. We expect analogous techniques would work to prove these results for uncountable groups, but require translating everything to the language of coarse structures. It is worth remarking that a countable group may have uncountably-many finite index subgroups.
@@
In order to avoid this problem, we assume a kind of compatibility between the family $(N_{i})$ and the particular extension. More precisely, we assume $(N_{i})$ separates $\Delta$ from $\Gamma$ (see LABEL:separable_subgroup_definition for a definition). We will see in LABEL:baumslag-solitar_groups_section that there are still significant applications, as all elementary amenable groups are built up from extensions which are compatible with sufficiently large collections. In fact, for our motivating example of the Baumslag-Solitar groups $\text{BS}(1,n)$, we will see that any sequence of finite index subgroups satisfies the hypotheses of our extension theorem. We show that box spaces of $BS(1,n)$ along filtrations have asymptotic dimension $2$ for all $n$, which in particular exhibits an infinite class of groups with exponential growth but finite-dimensional box spaces. In obtaining the lower bound for this result, we compute the asymptotic dimension of $BS(1,n)$, which by itself involves some interesting methods in geometric group theory. Specifically the notion of ends of groups allows one to prove that a torsion free group of asymptotic dimension $1$ is free (as we will see, this is an immediate consequence of results already in the literature). We also deal with some examples of more complicated groups which have similar presentations or are in some sense ‘generalized Baumslag-Solitar groups’.
More generally, it is natural to try to use our extension theorem along similar lines as in [finnsell2015asymptotic, Section 4], which attempts to relate the asymptotic dimension of box spaces of residually finite, elementary amenable groups to their Hirsch length. While the author encountered difficulty implementing this idea because of the additional assumptions in the extension theorem, we will still discuss how the content of this paper could be used to recover similar results. Specifically, one can probably use the Hurewicz theorem for the asymptotic dimension together with the ideas in section LABEL:applications to obtain an extension theorem for box spaces which doesn’t require the collection $(N_{i})$ to be a sequence; then use this to prove some special cases of the claims found in [finnsell2015asymptotic, Section 4].
@@
The following concept is originally due to Gromov, specifically [Gromov1991AsymptoticIO, Section 1.E]. There are many definitions, but we will use one throughout which makes the analogy with dynamic asymptotic dimension obvious:
@@
Let $(X,d)$ be a metric space. If $U\subset X$ is a subset, an $r$-chain in $U$ is a sequence $x_{0},\ldots,x_{n}$ of points in $U$ such that $d(x_{i},x_{i+1})\leq r$ for all $i$; and two points in $U$ are in the same $r$-component of $U$ if they are connected by an $r$-chain with all points contained in $U$. The subset $U$ is said to be $r$-connected if any two points in $U$ are connected by an $r$-chain in $U$. Finally, $\text{asdim}X\leq n$ means for all $r>0$ there is $M>0$ and a cover $\mathcal{U}=\{U_{0},\ldots,U_{n}\}$ of $X$ such that each $r$-component of each $U_{i}$ has diameter at most $M$. Such a cover is called a $(d,R,M)$-cover of $X$. The asymptotic dimension of $X$ is then the smallest integer $d$ such that $\text{asdim}X\leq d$, and is infinite if there is no such $d$.
@@
Many of the coarse spaces we consider will be graphs, considered as metric spaces by equipping their vertex set with the $l^{1}$-path metric. Finitely generated groups can be considered as metric spaces through the Cayley graph construction. More precisely, if $F\subset\Gamma$ is a finite generating set we denote by $C_{F}(\Gamma)$ the graph whose vertices are the elements of $\Gamma$ and which has an edge between $\gamma$ and $\delta$ if there is $f\in F$ such that $f\gamma=\delta$. For our purposes, we can always assume $F=F^{-1}$ so that we can think of $C_{F}(\Gamma)$ as an undirected graph. The $l^{1}$-path metric on the set of vertices is then the same as the word metric $d_{F}(\gamma,\delta)=|\gamma\delta^{-1}|_{F}$ (where $|\cdot|_{F}$ denotes the minimal length according to the generating set $F$). More generally, any countable group, if $\Gamma$ is a countable group, it has a right-invariant proper metric unique up to coarse equivalence [notes_on_prop_A, Lemma 2.3.3]. We will fix, from now on, some such metric on $\Gamma$ and note that the uniqueness up to coarse equivalence implies we can sensibly refer to $\text{asdim}\Gamma$.
@@
Throughout, $\Gamma\curvearrowright X$ and $\Lambda\curvearrowright Y$ will be actions on compact metric spaces by homeomorphisms.
@@
Let $\Gamma\curvearrowright X$ be an action on a compact Hausdorff space by homeomorphisms. The dynamic asymptotic dimension $\Gamma\curvearrowright X$, denoted $\text{DAD}(\Gamma\curvearrowright X)$, is the smallest integer $d$ such that for all finite $F\subset\Gamma$ there exists an open cover $\mathcal{U}=\{U_{0},\ldots,U_{d}\}$ of $X$ such that, for all $0\leq j\leq d$, the set
@@
The following terminology for DAD makes the analogy with asymptotic dimension clear, and will be critical to the proof of our main theorem. It is inspired by [Brodskiy2006AHT].
@@
Let $\mathcal{P}_{\text{fs}}(\Gamma)$ denote the collection of all finite, symmetric subsets of $\Gamma$ containing the identity. Let $X$ be a compact metrizable space, $\Gamma\curvearrowright X$ an action by homeomorphisms, $A\subset X$ a subset, and $F\in\mathcal{P}_{\text{fs}}(\Gamma)$. An $F$-chain in $A$ is a finite sequence $x_{0},\ldots,x_{n}$ of points in $A$ such that $x_{i+1}=f_{i+1}\cdot x_{i}$ with $f_{i+1}\in F$ ($0\leq i<n$). Two points $x,y\in A$ are in the same $F$-component of $A$ if they are connected by an $F$-chain in $A$ (this is an equivalence relation since $F$ is symmetric). We say $A$ is $F$-connected if any two points in $A$ can be connected by an $F$-chain. If $S\in\mathcal{P}_{\text{fs}}(\Gamma)$, we say $A\subset X$ is $S$-bounded if it is a subset of $S\cdot x$ for some $x\in X$. A cover of $A$ by $m+1$ sets which are open (respectively, closed) in the subspace topology on $A$ and whose $F$-components are $S$-bounded is called an open (respectively, closed) $(m,F,S)$-cover for $A$. Notice that in this case $S$ can always be taken to be $F^{r}$ for some $r>0$. In case we need to distinguish between different actions on $X$ (as we will in section LABEL:applications), we will refer to an (open/closed) $(m,F,S)$-cover of $A$ with respect to $\Gamma\curvearrowright X$.
An $n$-dimensional control function for $A$ (with respect to $\Gamma\curvearrowright X$) is a function $\mathcal{D}_{A}:\mathcal{P}_{\text{fs}}(\Gamma)\to\mathcal{P}_{\text{fs}}(\Gamma)$ such that for every $F\in\mathcal{P}_{\text{fs}}(\Gamma)$, there is an $(n,F,\mathcal{D}_{X}(F))$-cover for $A$ (again, with respect to $\Gamma\curvearrowright X$ if the action is ambiguous). Note that $\text{DAD}_{free}(\Gamma\curvearrowright X)\leq n$ iff there is an $n$-dimensional control function for $X$. We similarly define $\text{DAD}(A)$ to be the least integer $n\geq 0$ such that there exists an $n$-dimensional control function for $A$.
@@
Although the terminology in 2.6 relates specifically to the definition of DAD for free actions, we will use it even for actions which are not free. The Hurewicz-type theorem that we will prove in section 4 will be stated using the definition for free actions. Note, however, that by the definition for non-free actions, any action with infinite stabilizers will be infinite dimensional. Since the two definitions are equivalent for an action with uniformly-finite stabilizers, and actions with uniformly-finite stabilizers can easily be related to free actions by ignoring some torsion elements, we do not really lose anything by restricting our attention in this way. From now on we will only use the definition for free actions aside from a brief mention of the general definition in 3.
@@
Suppose $\mathcal{V}=\{V_{0},\ldots,V_{d}\}$ is an open cover of a totally disconnected, compact space $X$. Then there is a clopen cover $\mathcal{W}=\{W_{0},\ldots,W_{d}\}$ such that $W_{i}\subset V_{i}$ for all $0\leq i\leq d$. In the context of the previous definition, this implies if $\mathcal{V}$ is a $(d,F,S)$-cover, then $\mathcal{W}$ is a $(d,F,S)$-cover.
@@
Suppose $\Gamma\curvearrowright X$ is a free action. Then $\text{DAD}(\Gamma\curvearrowright X)\geq\text{asdim}\Gamma$. In fact, this inequality holds even if one modifies the definition of DAD to allow covers by Borel sets, or even arbitrary sets.
Proof.
Fix $r>0$. Let $F=B_{e}^{r}(\Gamma)$. Suppose $\text{DAD}(\Gamma\curvearrowright X)\leq d$. Then for some $S\in\mathcal{P}_{\text{fs}}(\Gamma)$, there is a $(d,F,S)$-cover, $\mathcal{V}$, for $\Gamma\curvearrowright X$. Let $\mathcal{O}(x)$ be the orbit of some $x\in X$. Identify $\mathcal{O}(x)$ with $\Gamma$ by $\gamma\cdot x\mapsto\gamma$. Then $\mathcal{V}$ gives rise to a cover $\mathcal{V}^{\prime}$ of $\Gamma$. Moreover, two points $y,y^{\prime}\in\mathcal{O}(x)$ are connected by an $F$ chain iff their corresponding elements in $\Gamma$ are connected by an $r$-chain. Therefore, the $r$-components of each $V\in\mathcal{V}^{\prime}$ are $S$-bounded, so $\mathcal{V}^{\prime}$ is a $(d,r,\text{diam}(S))$-cover for $C_{F}(\Gamma)$. ∎
@@
If $\{G_{n}\}_{n\in\mathbb{N}}$ is a countable collection of finite metric spaces, the coarse disjoint union $\sqcup_{n}G_{n}$ is the metric space with underlying set the disjoint union of the $G_{n}$ and metric $d(x,y):=d_{G_{n}}(x,y)$ if $x,y\in G_{n}$ and $d(x,y):=\text{diam}(G_{n})+\text{diam}(G_{m})$ if $x\in G_{n}$ and $y\in G_{m}$ with $m\neq n$. If $\Gamma$ is a group and $\{N_{i}\}$ is a collection of finite-index subgroups, we define $\Box_{(N_{i})}\Gamma:=\sqcup_{i}\Gamma/N_{i}$ where each $\Gamma/N_{i}$ has the metric induced from the right-invariant proper metric on $\Gamma$ we fixed earlier (i.e. the distance between cosets is the infimum over the distances between their different elements according to the original metric on $\Gamma$). This is called a box space of $\Gamma$. If the metric on $\Gamma$ is the word metric coming from some generating set and if the $N_{i}$ are normal, the induced metrics on quotients are the word metrics coming from the image of that generating set in the quotient.
@@
A filtration of $\Gamma$ is a collection $\{N_{i}\}$ of finite index subgroups of $\Gamma$ such that $\bigcap_{i}N_{i}=\{e\}$, the trivial subgroup. If $\Gamma$ is residually finite, the collection of all finite, index normal subgroups is then a filtration. Note that a countable, residually finite group has a filtration which is also a sequence, that is $N_{i}\supset N_{i+1}$.
@@
If $(G_{i})$ is a sequence of metric spaces, we say $\text{asdim}(G_{i})\leq d$ uniformly if for all $r>0$ there is an $R>0$ such that each $G_{i}$ admits a $(d,r,R)$-cover. We similarly say for a sequence $(\Gamma\curvearrowright X_{i})$ of actions on spaces that $\text{DAD}(\Gamma\curvearrowright X_{i})\leq d$ uniformly if for all $F\in\mathcal{P}_{\text{fs}}(\Gamma)$ there is $S\in\mathcal{P}_{\text{fs}}(\Gamma)$ such that there is a $(d,F,S)$-cover for each action. We can similarly define what it means for a general collection of metric spaces or actions to have asdim or DAD at most $d$ uniformly. In terms of control functions, this means having one control function which ‘works’ for all spaces or actions in the collection.
@@
$\text{asdim}\Box_{(N_{i})}\Gamma\leq d$ if and only if $\text{asdim}(\Gamma/N_{i})\leq d$ uniformly.
@@
For one inequality, notice that a $(d,r,R)$-cover for $\Box_{(N_{i})}\Gamma$ gives rise to a $(d,r,R)$-cover for each $\Gamma/N_{i}$ (just take the intersection).
@@
The following lemma shows that the dimension of groups and box spaces behaves nicely with respect to direct limits. While we will not really use it, it is relevant to the discussion of elementary-amenable groups in LABEL:baumslag-solitar_groups_section and moreover does not seem to be written down anywhere.
@@
Suppose $G=\bigcup_{i}G_{i}$ is an increasing union of groups and $X=\Box_{(H_{j})}G=\bigcup_{i}X_{i}$ is an increasing union of box spaces where $X_{i}=\Box_{(H_{j}\cap G_{i})}G_{i}$ is a box space of $G_{i}$ (here the metrics on $X$ and $G_{i}$ are induced from the right invariant, proper metric on $G$, and the metric on $X_{i}$ is constructed using the metric on $G_{i}$). Then $\text{asdim}G=\sup_{i}\text{asdim}G_{i}$ and $\text{asdim}X=\sup_{i}\text{asdim}X_{i}$.

### v6

@@ 1 Introduction
More recently, Hurewicz-type theorems were developed for Gromov’s asymptotic dimension in [asymptotichurewicztheorem] and later in [Brodskiy2006AHT]. This paper uses similar methods as in [Brodskiy2006AHT], together with some original ideas, to prove a Hurewicz-type theorem for the dynamic asymptotic dimension (DAD) of Guentner, Willett, and Yu. This dimension theory was introduced in [dasdimGWY] as a dynamical analogue of the asymptotic dimension. It is related to conditions used in work on the Farrell-Jones conjecture on manifold topology [bartels] [bartels_et_al], as well as to the homology of group actions [bonicke2021dynamic] and the nuclear dimension of transformation group $C^{*}$-algebras [dasdimGWY, Theorem 8.6]. It has also been applied to the hyperfiniteness problem for orbit equivalence relations [Conley2020BorelAD].
@@
Suppose $1\to\Delta\to\Gamma\to\Lambda\to 1$ is an exact sequence of countable groups and that $\Gamma$ is residually finite. Suppose $(N_{i})$ is a countable collection of finite index normal subgroups of $\Gamma$ which separates $\Delta/(\cap_{i}(N_{i}\cap\Delta))$ from $\Gamma/(\cap_{i}N_{i})$ (see 6.3). This collection induces collections of finite index normal subgroups and hence odometers of $\Delta$ and $\Lambda$. Then if $\text{DAD}_{free}(\Delta\curvearrowright\widehat{\Delta}_{(N_{i}\cap\Delta)})=m$ and $\text{DAD}_{free}(\Lambda\curvearrowright\widehat{\Lambda}_{(N_{i}/\Delta\cap N_{i})})=n$, $\text{DAD}_{free}(\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})})\leq m+n$.
This extension theorem can further be applied to coarse geometry, specifically the study of box spaces of groups. These are spaces constructed out of finite quotients of a group. Just as the asymptotic dimension of a group itself can have interesting group-theoretic applications (see for instance [MR2431020]), there are likewise many correspondences between geometric properties of box spaces and group-theoretic properties. For example, coarse equivalence of box spaces encodes virtual isomorphism of groups [MR4091056], and amenability of groups is equivalent to property A of their box spaces [Roe2003LecturesOC, Proposition 11.39]. Similar results also exist for a-T-menablity [https://doi.org/10.1112/blms/bdt045].
Box spaces provided some of the first examples of spaces which do not coarsely embed in Hilbert space [Roe2003LecturesOC, Remark 11.27] and it has been shown that a group with a finite dimensional box space must be amenable [Roe2003LecturesOC, Remark 11.38 and Proposition 11.39]. Just as actions by finite dimensional, amenable groups are thought to be finite dimensional, it is thought that a partial converse to this latter fact should also hold; that is, an amenable group with finite asymptotic dimension ought to have finite dimensional box spaces. Notwithstanding, there has been similarly little progress on this aside from the case of virtually nilpotent groups [boxspacesDT].
@@
Let $\Gamma$ be a countable group and $(N_{i})$ a sequence of finite index subgroups of $\Gamma$. Suppose $\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})}$ is the action by left multiplication of $\Gamma$ on its profinite completion along $(N_{i})$. Then $\text{DAD}_{free}(\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})})=\operatorname{\text{asdim}}\Box_{(N_{i})}\Gamma$.
@@
Suppose $1\to\Delta\to\Gamma\xrightarrow{\pi}\Lambda\to 1$ is an exact sequence of countable groups and $(N_{i})$ is a sequence of finite index normal subgroups of $\Gamma$ which separates $\Delta/(\cap_{i}(N_{i}\cap\Delta))$ from $\Gamma/(\cap_{i}N_{i})$. This sequence induces sequences and hence box spaces of $\Delta$ and $\Lambda$. Then if $\operatorname{\text{asdim}}\Box_{(N_{i}\cap\Delta)}\Delta=m$ and $\operatorname{\text{asdim}}\Box_{(\pi(N_{i}))}\Lambda=n$, $\operatorname{\text{asdim}}\Box_{(N_{i})}\Gamma\leq m+n$.
We have therefore established a special case of the (unproved) statement [finnsell2015asymptotic, Theorem A]. Our extra assumptions are that $\Gamma$ is countable, that the $N_{i}$ are normal, and that $(N_{i})$ is a sequence separating $\Delta$ from $\Gamma$. In [finnsell2015asymptotic, Theorem A], the family of subgroups are only assumed to be semiconjugacy separating, which is weaker than assuming $(N_{i})$ to be a filtration by normal subgroups. We expect analogous techniques would work to prove these results for uncountable groups, but require translating everything to the language of coarse structures. It is worth remarking that a countable group may have uncountably-many finite index subgroups.
@@
In order to avoid this problem, we assume a kind of compatibility between the family $(N_{i})$ and the particular extension. More precisely, we assume $(N_{i})$ separates $\Delta$ from $\Gamma$ (see 6.3 for a definition). We will see in 7 that there are still significant applications, as all elementary amenable groups are built up from extensions which are compatible with sufficiently large collections. In fact, for our motivating example of the Baumslag-Solitar groups $\text{BS}(1,n)$, we will see that any sequence of finite index subgroups satisfies the hypotheses of our extension theorem. We show that box spaces of $BS(1,n)$ along filtrations have asymptotic dimension $2$ for all $n$, which in particular exhibits an infinite class of groups with exponential growth but finite-dimensional box spaces. In obtaining the lower bound for this result, we compute the asymptotic dimension of $BS(1,n)$, which by itself involves some interesting methods in geometric group theory, specifically the notion of ends of groups, which allows one to prove that a torsion free group of asymptotic dimension $1$ is free. As we will see, this is an immediate consequence of results already in the literature. We also deal with some examples of more complicated groups which have similar presentations or are in some sense ‘generalized Baumslag-Solitar groups’.
More generally, it is natural to try to use our extension theorem along similar lines as in [finnsell2015asymptotic, Section 4], which attempts to relate the asymptotic dimension of box spaces of residually finite, elementary amenable groups to their Hirsch length. While the author encountered difficulty implementing this idea because of the additional assumptions in the extension theorem, we will still discuss how the content of this paper could be used to recover similar results. Specifically, one can probably use the Hurewicz theorem for the asymptotic dimension together with the ideas in section 6 to obtain an extension theorem for box spaces which doesn’t require the collection $(N_{i})$ to be a sequence; then use this to prove some special cases of the claims found in [finnsell2015asymptotic, Section 4].
@@
The following concept is originally due to Gromov, specifically [Gromov1991AsymptoticIO, Section 1.E]. There are many definitions, but we will use one throughout which makes the analogy with the dynamic asymptotic dimension obvious:
@@
Let $(X,d)$ be a metric space. If $U\subset X$ is a subset, an $r$-chain in $U$ is a sequence $x_{0},\ldots,x_{n}$ of points in $U$ such that $d(x_{i},x_{i+1})\leq r$ for all $i$; and two points in $U$ are in the same $r$-component of $U$ if they are connected by an $r$-chain in $U$. The subset $U$ is said to be $r$-connected if any two points in $U$ are connected by an $r$-chain in $U$. Finally, $\operatorname{\text{asdim}}X\leq n$ means for all $r>0$ there is $M>0$ and a cover $\mathcal{U}=\{U_{0},\ldots,U_{n}\}$ of $X$ such that each $r$-component of each $U_{i}$ has diameter at most $M$. Such a cover is called a $(d,R,M)$-cover of $X$. The asymptotic dimension of $X$ is then the smallest integer $d$ such that $\operatorname{\text{asdim}}X\leq d$ and is infinite if no such $d$ exists.
@@
Many of the coarse spaces we consider will be graphs, considered as metric spaces by equipping their vertex set with the $l^{1}$-path metric. Finitely generated groups can be considered as metric spaces through the Cayley graph construction. More precisely, if $F\subset\Gamma$ is a finite generating set we denote by $C_{F}(\Gamma)$ the graph whose vertices are the elements of $\Gamma$ and which has an edge between $\gamma$ and $\delta$ if there is $f\in F$ such that $f\gamma=\delta$. For our purposes, we can always assume $F=F^{-1}$ so that we can think of $C_{F}(\Gamma)$ as an undirected graph. The $l^{1}$-path metric on the set of vertices is then the same as the word metric $d_{F}(\gamma,\delta)=|\gamma\delta^{-1}|_{F}$ (where $|\cdot|_{F}$ denotes the minimal length according to the generating set $F$). More generally, if $\Gamma$ is a countable group, it admits a right-invariant proper metric unique up to coarse equivalence [notes_on_prop_A, Lemma 2.3.3]. We will fix, from now on, some such metric on $\Gamma$ and note that its uniqueness up to coarse equivalence implies we can sensibly refer to $\operatorname{\text{asdim}}\Gamma$.
@@
Throughout, $\Gamma\curvearrowright X$ and $\Lambda\curvearrowright Y$ will be actions of countable groups on compact metric spaces by homeomorphisms.
@@
Let $\Gamma\curvearrowright X$ be an action on a compact Hausdorff space by homeomorphisms. The dynamic asymptotic dimension of $\Gamma\curvearrowright X$, denoted $\text{DAD}(\Gamma\curvearrowright X)$, is the smallest integer $d$ such that for all finite $F\subset\Gamma$ there exists an open cover $\mathcal{U}=\{U_{0},\ldots,U_{d}\}$ of $X$ such that, for all $0\leq j\leq d$, the set
@@
The following terminology makes the analogy with the asymptotic dimension clear, and will be critical to the proof of our main theorem. It is inspired by [Brodskiy2006AHT].
@@
Let $\mathcal{P}_{\text{fs}}(\Gamma)$ denote the collection of all finite, symmetric subsets of $\Gamma$ containing the identity. Let $X$ be a compact metrizable space, $\Gamma\curvearrowright X$ an action by homeomorphisms, $A\subset X$ a subset, and $F\in\mathcal{P}_{\text{fs}}(\Gamma)$. An $F$-chain in $A$ is a finite sequence $x_{0},\ldots,x_{n}$ of points in $A$ such that $x_{i+1}=f_{i+1}\cdot x_{i}$ with $f_{i+1}\in F$ ($0\leq i<n$). Two points $x,y\in A$ are in the same $F$-component of $A$ if they are connected by an $F$-chain in $A$ (this is an equivalence relation since $F$ is symmetric). We say $A$ is $F$-connected if any two points in $A$ can be connected by an $F$-chain. If $S\in\mathcal{P}_{\text{fs}}(\Gamma)$, we say $A\subset X$ is $S$-bounded if it is a subset of $S\cdot x$ for some $x\in X$. A cover of $A$ by $m+1$ sets which are open (respectively, closed) in the subspace topology on $A$ and whose $F$-components are $S$-bounded is called an open (respectively, closed) $(m,F,S)$-cover of $A$. Notice that in this case $S$ can always be taken to be $F^{r}$ for some $r>0$. In case we need to distinguish between different actions on $X$ (as we will in section 6), we will refer to an (open/closed) $(m,F,S)$-cover of $A$ with respect to $\Gamma\curvearrowright X$.
An $n$-dimensional control function for $A$ (with respect to $\Gamma\curvearrowright X$) is a function $\mathcal{D}_{A}:\mathcal{P}_{\text{fs}}(\Gamma)\to\mathcal{P}_{\text{fs}}(\Gamma)$ such that for every $F\in\mathcal{P}_{\text{fs}}(\Gamma)$, there is an open $(n,F,\mathcal{D}_{X}(F))$-cover of $A$ (again, with respect to $\Gamma\curvearrowright X$ if the action is ambiguous). Note that $\text{DAD}_{free}(\Gamma\curvearrowright X)\leq n$ iff there is an $n$-dimensional control function for $X$. We similarly define $\text{DAD}(A)$ to be the least integer $n\geq 0$ such that there exists an $n$-dimensional control function for $A$.
@@
Although the terminology in 2.6 relates specifically to the definition of DAD for free actions, we will use it even for actions which are not free. The Hurewicz-type theorem that we will prove in section 4 will be stated using the definition for free actions. Note, however, that by the definition for non-free actions, any action with infinite stabilizers will be infinite dimensional. Since the two definitions are equivalent for an action with uniformly-finite stabilizers, and actions with uniformly-finite stabilizers can easily be related to free actions by ignoring some torsion elements, we do not really lose anything by restricting our attention in this way. From now on we will only use the definition for free actions aside from a brief mention of the general definition in section 3.
@@
Suppose $\mathcal{V}=\{V_{0},\ldots,V_{d}\}$ is an open cover of a totally disconnected, compact space $X$. Then there is a clopen cover $\mathcal{W}=\{W_{0},\ldots,W_{d}\}$ such that $W_{i}\subset V_{i}$ for all $0\leq i\leq d$. In the context of the previous definition, this implies $\mathcal{W}$ is a $(d,F,S)$-cover if $\mathcal{V}$ is a $(d,F,S)$-cover.
@@
The following lemma appeared first in [toprigiditydad, Lemma 2.9], and we refer there for a proof.
@@
Suppose $\Gamma\curvearrowright X$ is a free action. Then $\text{DAD}(\Gamma\curvearrowright X)\geq\operatorname{\text{asdim}}\Gamma$. In fact, this inequality holds even if one modifies the definition of DAD to allow covers by Borel sets, or even arbitrary sets. ∎
@@
If $\{G_{n}\}_{n\in\mathbb{N}}$ is a countable collection of finite metric spaces, the coarse disjoint union $\sqcup_{n}G_{n}$ is the metric space with underlying set the disjoint union of the $G_{n}$ and metric $d(x,y):=d_{G_{n}}(x,y)$ if $x,y\in G_{n}$ and $d(x,y):=\text{diam}(G_{n})+\text{diam}(G_{m})$ if $x\in G_{n}$ and $y\in G_{m}$ with $m\neq n$. If $\Gamma$ is a group and $\{N_{i}\}$ is a collection of finite-index subgroups, we define $\Box_{(N_{i})}\Gamma:=\sqcup_{i}\Gamma/N_{i}$ where each $\Gamma/N_{i}$ has the metric induced from the right-invariant proper metric on $\Gamma$ we fixed earlier (i.e. the distance between cosets is the infimum over the distances between their different elements according to the original metric on $\Gamma$). This is called a box space of $\Gamma$. If the metric on $\Gamma$ is the word metric coming from some generating set, and if the $N_{i}$ are normal, the induced metrics on quotients are the word metrics coming from the image of that generating set in the quotient.
@@
A filtration of $\Gamma$ is a collection $\{N_{i}\}$ of finite index subgroups of $\Gamma$ such that $\bigcap_{i}N_{i}=\{e\}$, the trivial subgroup. If $\Gamma$ is residually finite, the collection of all finite index normal subgroups is then a filtration. Note that a countable, residually finite group has a filtration which is also a sequence, that is $N_{i}\supset N_{i+1}$.
@@
If $(G_{i})$ is a sequence of metric spaces, we say $\operatorname{\text{asdim}}(G_{i})\leq d$ uniformly if for all $r>0$ there is an $R>0$ such that each $G_{i}$ admits a $(d,r,R)$-cover. We similarly say for a sequence $(\Gamma\curvearrowright X_{i})$ of actions on spaces that $\text{DAD}(\Gamma\curvearrowright X_{i})\leq d$ uniformly if for all $F\in\mathcal{P}_{\text{fs}}(\Gamma)$ there is $S\in\mathcal{P}_{\text{fs}}(\Gamma)$ such that there is an open $(d,F,S)$-cover of each $X_{i}$. We can similarly define what it means for a general collection of metric spaces or actions to have $\operatorname{\text{asdim}}$ or DAD at most $d$ uniformly. In terms of control functions, this means having one control function which ‘works’ for all spaces or actions in the collection.
@@
$\operatorname{\text{asdim}}\Box_{(N_{i})}\Gamma\leq d$ if and only if $\operatorname{\text{asdim}}(\Gamma/N_{i})\leq d$ uniformly.
@@
For one inequality, notice that a $(d,r,R)$-cover of $\Box_{(N_{i})}\Gamma$ gives rise to a $(d,r,R)$-cover of each $\Gamma/N_{i}$ (just take the intersection).
@@
The following lemma shows that the dimension of groups and box spaces behaves nicely with respect to direct limits. While we will not really use it, it is relevant to the discussion of elementary-amenable groups in 7 and does not seem to be written down anywhere.
@@
Suppose $G=\bigcup_{i}G_{i}$ is an increasing union of groups and $X=\Box_{(H_{j})}G=\bigcup_{i}X_{i}$ is an increasing union of box spaces where $X_{i}=\Box_{(H_{j}\cap G_{i})}G_{i}$ is a box space of $G_{i}$ (here the metrics on $X$ and $G_{i}$ are induced from the right invariant, proper metric on $G$, and the metric on $X_{i}$ is constructed using the metric on $G_{i}$). Then $\operatorname{\text{asdim}}G=\sup_{i}\operatorname{\text{asdim}}G_{i}$ and $\operatorname{\text{asdim}}X=\sup_{i}\operatorname{\text{asdim}}X_{i}$.